Specifically, you are to prepare a graph of Absorbance versus [Co2+] for the following solutions -- prepared by diluting the specified volumes of 0.2500 M Co2+ stock solution to 25.00 mL each:

volume Absorbance
0mL 0
2.50mL .165
5.00mL .320
7.50 mL .485
10.0mL .610
12.0mL .721

What is the slope?

For an unknown solution whose absorbance was 0.330, what would be the corresponding value of [Co2+]?
Answer #1

Let's first calculate concentration of Co2+ in all the solutions:
Formula: M1V1 = M2V2
Solution-2
M1 = 0.2500 M, V1 = 2.50mL, V2 = 25mL
So M2 = (0.25*2.5)/25 = 0.025 M
Solution-3
M1 = 0.2500 M, V1 = 5.00mL, V2 = 25mL
So M2 = (0.25*5)/25 = 0.05 M
Solution-4
M1 = 0.2500 M, V1 = 7.50mL, V2 = 25mL
So M2 = (0.25*7.5)/25 = 0.075 M
Solution-5
M1 = 0.2500 M, V1 = 10mL, V2 = 25mL
So M2 = (0.25*10)/25 = 0.1 M
Solution-6
M1 = 0.2500 M, V1 = 12mL, V2 = 25mL
So M2 = (0.25*12)/25 = 0.12 M

A B C D E F G H I J 1 Concentration Absorbance The equation of best fit line is: o 0 Y = 6.012*X +0.012 0.025 0.165 0.05 0.32